OpenStudy (anonymous):

Find the amount of an annuity that consists of 29 monthly payments of $600 each into an account that pays 3% interest per year, compounded monthly. (Round your answer to the nearest cent.) I got 27,131.35 help its wrong -_-!

OpenStudy (anonymous):

so pissed off what is an easier way to do it in a calculator !?

OpenStudy (mathmate):

Try using i=0.03/12=0.0012 n=29 P=payment per period Amount accumulated =(P(1+i)^n)/i I get about $18023